OpenStudy (anonymous):

Anyone who knows derivatives?I need help because i'm having a hard time understanding derivatives having trigo functions and derivative of square root.

OpenStudy (anonymous):

how to solve this \[\sqrt{x+1}\]

OpenStudy (anonymous):

Use the chain rule

OpenStudy (anonymous):

u r wrong........................ zhanga

OpenStudy (anonymous):

Crap I am.

OpenStudy (anonymous):

but anyway, use chain rule.

OpenStudy (anonymous):

zhanga was right about applying chain rule but wrong with the answer hahahaha. let u = x + 1 du/dx = 1 dy/dx = dy/du * du/dx dy/dx(sqrt(u)) = 1 / (2(sqrt(x + 1))) * 1 therefore the derivative of sqrt(x+1) is 1 / (2sqrt(x + 1))

OpenStudy (anonymous):

\[y=\sqrt{x+1}\]Let u = x+1 Therefore, du/dx = 1 and dy/du = 1/2u^-1/2 To find dy/dx you have to multiply dy/du and du/dx, Therefore, dy/dx = 1/2(x+1)^-1/2

OpenStudy (anonymous):

but how to solve it using the 4 step rule?

OpenStudy (anonymous):

come to twiddla

OpenStudy (anonymous):

http://www.twiddla.com/584540

OpenStudy (anonymous):

easy way: sqrt(x+1) (x+1)^(1/2) The step above is changing from square root form to exponent form. Something to the one half power means "square root it". So now we use chain rule: 1/2(x+1)^(-1/2) is the first step. All we did was power rule: drop the power in front, and subtract 1/2 by "one", which would be 2/2 in fraction form, leaving us with negative one half. But we're not done yet. We have to find the derivative of the inside junk (x+1). That would be 1+0 (x turns into 1, and the +1 just gets turned into 0. So: 1/2 ((x+1) ^(1/2))*1 or... \[\frac{1}{2\sqrt{x+1}}\]

OpenStudy (anonymous):

*So: 1/2 ((x+1) ^(-1/2))*1 FOrgot to put my negative sign in my original "so: blah blah" thing. The final answer is still the same, though.

OpenStudy (anonymous):

hey sheena are you done with diff. calculus?